A particle carrying charge q enters a magnetic field B with velocity v at an angle
Θ≠0≠90° . Show that it follows a helical path. Determine the radius of the helix.
1
Expert's answer
2020-02-21T09:59:47-0500

As per the question,

Charge of the particle =q

Magnetic field =B

Velocity =v

Let the angle between the magnetic field and the velocity is "\theta"

So,


velocity along the tangential direction to the helix =v"\sin\theta"

now, centripetal force =magnetic force

"\dfrac{m(v\sin\theta)^2}{R}=qvB\sin\theta"

"\Rightarrow R=\dfrac{qB}{mv\sin\theta}"

here R is the radius of the helix.

The horizontal component of the velocity is responsible for the motion in the horizontal direction of the helix.